Ashley brought some biscuits for morning tea at work. Her colleagues ate 79of them , then her boss came by and ate two biscuits. When Ashley came to eat some, there were only four left. How many biscuits did she bring for morning tea?
Answer by Boreal(15235)

x=brought
x-79-2=4
x=79+2+4=85 biscuits